OpenStudy (anonymous):

Find all solutions: x2 = 36 A) ±3 2 B) 6 C) ±6 D) 36

OpenStudy (jiteshmeghwal9):

take square root both sides :)

OpenStudy (anonymous):

\[x^2 = 36\] \[x = \pm 6\]

OpenStudy (hba):

Well. What about putting a square root on both sides, Lemme show you an example, \[\huge\ x^2=25 \] \[\sqrt{x^2}=\sqrt{25}\] \[x= \pm 5\]

OpenStudy (jiteshmeghwal9):

\[\sqrt{x^2}=\sqrt{36}\]

OpenStudy (anonymous):

the square root of 36 is 6

OpenStudy (hba):

Best of luck with this question @Pookiieluvzyhu

OpenStudy (hba):

Yes what about the sign ?

OpenStudy (hba):

@Pookiieluvzyhu See my example.

OpenStudy (jiteshmeghwal9):

sqrt always gets the sign of \(\pm\) since sqrt{25}=5 if we took (5)^2=25 & (-5)^2 too 25

OpenStudy (anonymous):

okay thanks for the example.but i got a question what does that plus sign over the line stand for

Parth (parthkohli):

Observe:\[6^2 =36\]\[(-6)^2 = 36\]

OpenStudy (anonymous):

kkk i got it now